Looking through photos from 03-09-12 I did not see any soot or charing around the downstream side of the Septa, where there is some today. The pictures are not the best, but seems it would have been evident. So, looking through the Archiver I found the serious problem became evident on the 14th after the trip at 18:10, noted in the halog Entry number 366496. It seems the Right Septum came back up with a short somewhere. The power supply is regulating on current control, it will output a fixed current. The hall probe in the rear of the Septum should read the appropriate field for the output current, if all the current is going through the coils in the correct path. At the left of the attached picture we can see that when the Right Septum current was restored, the hall probe(blue trace) was reading a lower field value than prior to the trip. The current in the Right Septum was not following the correct path. We can see that this continues, with some variations, till noon on the 16th. We can see in Entry number 366740 that the rate in the RHRS was back to "normal" at around 17:40. We do not have a readback of the output voltage, but we most likely would have seen the output voltage had risen to drive the current through the charred splice on the top of the Right Septum. Eventually something got to be too much for the supply and it tripped off this morning.